determine the validity of the statement below, and illustrate how you have arroved at your inference.
i. y-2=3(x-2)
ii. y=X-4

Answer by venugopalramana(3286)

determine the validity of the statement below, and illustrate how you have arroved at your inference.
i. y-2=3(x-2).....Y-2=3X-6..OR....3X-Y=4.............1
ii. y=X-4.........................X-Y =4.................2
EQN.1 - EQN.2...
3X-Y-X+Y=4-4=0
2X=0....OR....X=0
HENCE Y=X-4=0-4=-4
HRNCE THE STATEMENTS ARE TRUE FOR X=0 AND Y=-4